# Approaching Stable Quark Matter

### v1

@@ Abstract
The determination of whether the ground state of baryon matter in Quantum Chromodynamics (QCD) is the ordinary nucleus or a quark matter state remains a long-standing question in physics. A critical parameter in this investigation is the bag parameter $B$, which quantifies the QCD vacuum energy and can be computed using nonperturbative methods such as Lattice QCD (LQCD). By combining the equation of state derived from perturbative QCD (pQCD) with the bag parameter to fit the LQCD-simulated data for isospin-dense matter, we address the stability of quark matter within the LQCD+pQCD framework. Our findings suggest that the current data imposes an upper bound on $B^{1/4}\lesssim 160$ MeV, approaching a conclusive statement on quark matter stability. Additionally, we derive more general thermodynamic bounds on the quark matter energy-per-baryon and $B$, which, while weaker, provide complementary insights.
@@ 1 Introduction
It has long been thought that the global ground state of the baryon matter in quantum chromodynamics (QCD) should be the hadron state, while the quark matter state could only be accessible at environments of high temperature (such as the early universe or ion colliders) or high density (such as the core of a massive neutron star). However, the possible form of quark matter existence as a solitonic “bag” in the zero-pressure vacuum, also known as the “quark nugget”, has long ago been hypothesized and studied extensively by physicists such as Bodmer [1], Lee and collaborators [2, 3, 4, 5], and Witten [6], and still receives wide attention today. In contrast to the possible quark matter at the core of a neutron star, whose stability is supported by the external pressure, the stability of a quark nugget relies entirely on the energy configuration of the matter itself. Usually, a quark matter state is considered stable if its energy per baryon $\epsilon/n_{\text{B}}$ is less than around $930$ MeV, the averaged mass per nucleon of the iron element. So far whether this is true still remains an open question. The question of quark matter stability is of major importance since the implied stable quark nuggets could potentially extend the periodic table by introducing “exotic nuclei” with very large atomic and atomic mass numbers. Furthermore, stable quark nuggets, if produced in the early universe, could serve as a compelling dark matter candidate [6].
To investigate this question, one can start from the high-density limit, within which the thermodynamic equation of state (EOS) of the quark matter phase can be calculated based on perturbative QCD (pQCD) [7, 8, 9]. For instance, Ref. [10] has performed such calculations up to $\mathcal{O}(\alpha_{s}^{2})$ including the strange quark mass, which we mainly follow in this work. The pressure from pQCD calculations is a positive function of the baryon chemical potential $\mu_{\text{B}}$. The additional formation of color superconducting (CS) phase from Bardeen-Cooper-Schrieffer condensates provides also positive but negligible pressure to the quark matter system in the perturbative regime. To achieve a quark matter state with zero external pressure, an additional negative contribution to the total pressure is required. This contribution could arise from the vacuum energy difference between the confined (chiral-symmetry breaking) and deconfined (chiral-symmetry restored) QCD phases, which is nonperturbative in nature.
@@
$\hskip 85.35826ptB\equiv\frac{1}{4}\,\langle\Theta^{\mu}_{\;\mu}\rangle\leavevmode\nobreak\ ,\qquad\mbox{with}\quad\Theta^{\mu}_{\;\mu}=\frac{\beta(g)}{2g}G_{a}^{\mu\nu}G^{a}_{\mu\nu}+\sum_{q}\,m_{q}[1+\gamma_{m}(g)]\overline{q}q\leavevmode\nobreak\ .$ (1.1)
Here, $\Theta^{\mu}_{\;\mu}$ is the trace anomaly that depends on the renormalized fields and the QCD gauge coupling $g$ [13, 14]; $\beta(g)$ is the beta-function of the gauge coupling and $\gamma_{m}(g)$ is the quark mass anomalous dimension. The vacuum expectation value of the trace anomaly operator is defined at zero temperature and density.
Given that it requires a nonperturbative method to determine the bag parameter, lattice QCD (LQCD) is by far the unique tool for this task. In order to explore the transition from the hadron phase to the quark matter phase, one can either probe along the temperature direction or along the chemical potential direction (see for example Ref. [15] for a review on the QCD phase diagram). In this work, we focus on the latter and will comment on the former in Section 7. It is well known that simulating a system at finite $\mu_{\text{B}}$ (LQCDB) suffers from the fermion sign problem [16], making it challenging to extend the study to a high $\mu_{\text{B}}$. On the other hand, the isospin-dense matter system (LQCDI) is free of this problem (see Section 4), and thus can be studied up to a high isospin chemical potential $\mu_{\text{I}}$ to explore the deconfined quark matter phase. Recently, Refs. [17, 18] have performed LQCD simulations on a system of 6144 pions up to $\mu_{\text{I}}\sim 3.2$ GeV at zero temperature, in which they have identified the first-order phase transition to a Bose-Einstein condensate phase and a crossover to the CS phase [19, 20] through measuring the pressure $p$, energy density $\epsilon$, and speed of sound $c_{s}$ of the system. Based on the next-to-leading-order (NLO) calculations of the CS contribution to the EOS derived in Ref. [21], Ref. [22] has performed a matching between the LQCDI data and the pQCD+CS calculations to estimate the CS gap parameter. In this work, we take into account the $B$ parameter together with pQCD+CS to analyze the LQCDI data. This will not only provide a more complete description of the system, but also potentially examine the stability of quark matter.
@@ 2 Phases of baryon matter
$-\frac{\Omega}{V}=p=\mu_{\text{B}}\,n_{\text{B}}-\epsilon\leavevmode\nobreak\ .$ (2.1)
@@
Figure 1: Left panel: A schematic plot of the $p(\mu_{\text{B}})$ curves of the hadron matter phase and the quark matter phase at zero temperature, when there is only one intersection at $\mu_{\text{B}}=\mu_{\rm t}$ (green square point) between the two phases. The blue circle point at $\mu_{\text{B}}=\mu_{\rm h}$ with $p=0$ gives the energy per baryon of the (free) hadron matter as $\mu_{\rm h}$, while the red circle point at $\mu_{\text{B}}=\mu_{\rm q}$ gives that of the quark matter. The lower endpoint of the quark matter phase at $\mu_{\text{B}}=\mu_{0}$ and $p=-B$ has $n_{\text{B}}=0$. The globally stable baryon matter phase (green solid line) follows the trajectory of the maximum $p$ at any given $\mu_{\text{B}}$. Right panel: The same as the left panel but with two intersections at $\mu_{\text{B}}=\mu_{{\rm t}_{1}}$ and $\mu_{{\rm t}_{2}}$ between the two phases.
@@
The ordinary hadron matter phase (blue dot-dashed line) starts from a nonzero $\mu_{\text{B}}=\mu_{\rm h}$ (bounded from below by the baryon mass), where both the pressure and the number density are zero. As for the quark matter phase (red dashed line), the starting chemical potential $\mu_{0}$ could lie at a lower value, where $n_{\text{B}}$ is zero but the pressure is non-positive, $p=-B$ with $B\geq 0$. As $\mu_{\text{B}}$ increases, there will be a zero-pressure point at $\mu_{\text{B}}=\mu_{\rm q}$ where the free quark matter state can exist with its energy per baryon given by $\epsilon/n_{\text{B}}=\mu_{\rm q}$ as derived from Eq. (2.1).
@@
The $p(\mu_{\text{B}})$ curves of the two phases are expected to have at least one intersection. In the high $\mu_{\text{B}}$ limit, one anticipates the quark matter phase (including the CS phase) to be the ultimate phase, and the hadron matter phase starting from $\mu_{\text{B}}=\mu_{\rm h}$ will eventually flow to the quark matter phase. In general, the $p(\mu_{\text{B}})$ curves of the two phases can intersect multiple times. For simplicity, we will consider the situation with only one intersection at $\mu_{\rm t}$ (shown in the left panel of Figure 1) and that with two at $\mu_{{\rm t}_{1,2}}$ (shown in the right panel of Figure 1). Using the Maxwell construction to patch the curves of the two different phase  [^1] Alternatively, the Gibbs construction further considers the electric charge chemical potential of the thermodynamic system., both the chemical potentials and pressures of the two phases should be continuous at the transition boundaries, i.e., $\mu_{1}=\mu_{2}=\mu_{\rm t}$ and $p_{1}(\mu_{\rm t})=p_{2}(\mu_{\rm t})$. For the case with one intersection in the left panel, one inevitably has $\mu_{\rm q}>\mu_{\rm h}$ or $\epsilon/n_{\text{B}}|_{\rm q}>\epsilon/n_{\text{B}}|_{\rm h}$, indicating an unstable quark matter state. For a fixed value of $\mu_{\rm t}$, $\mu_{\rm q}$ will increase along with increasing $B$, making the quark matter state more unstable. Conversely, $\mu_{\rm q}<\mu_{\rm h}$ for the case of two intersections in the right panel, resulting in a more stable quark matter state compared to the ordinary hadron matter state.
@@ 3 Quark matter stability from pQCD calculations
At zero temperature and a high enough chemical potential, the grand potential or the pressure of the quark matter can be calculated perturbatively by treating the QCD coupling $\alpha_{s}$ as an expansion parameter [7, 8, 9, 10] and including the single quark loops, two-gluon irreducible vacuum diagrams, and the plasmon ring sum (see Ref. [10] for details). In this section, we follow Ref. [10] to calculate the thermodynamic properties of the quark matter system up to $\mathcal{O}(\alpha_{s}^{2})$. Together with the unknown bag parameter $B$, which can only be calculated nonperturbatively, we derive the stability conditions for the quark matter.
@@
$X=\begin{cases}\dfrac{\mu_{\text{R}}}{\frac{2}{N_{c}}\left(\mu_{u}+\mu_{d}+\mu_{s}\right)}&\leavevmode\nobreak\ \text{[2+1-flavor]}\\[12.91663pt] \dfrac{\mu_{\text{R}}}{\frac{2}{N_{c}}\left(\mu_{u}+2\mu_{d}\right)}&\leavevmode\nobreak\ \text{[2-flavor]}\\[12.91663pt] \dfrac{\mu_{\text{R}}}{\mu_{u}-\mu_{d}}&\leavevmode\nobreak\ \text{[isospin-dense]}\end{cases}\leavevmode\nobreak\ ,$ (3.1)
@@
Figure 2: Left panel: The $p(\mu_{\text{B}},X)=p_{\rm pQCD}+p_{\rm CS}$ distributions of the benchmarks with $X=1.6$ and $2.2$ for the 2+1- and 2-flavor systems. The black dashed vertical line stands for $\mu_{\text{B}}=930$ MeV, whose intersections with the $p$ distributions imply the upper bounds on the bag parameter $B$ that allow stable quark matter. Right panel: The parameter region (white) of stable quark matter with trustable pQCD calculations. The regions above the blue solid and orange dashed lines exclude the existence of stable 2+1- and 2-flavor quark matter, respectively. The pQCD calculation breaks down for $\mu_{\text{B}}\leq 930$ MeV in the shaded gray region of $X\leq 1.42$ (see Appendix A for details) and thus cannot give conclusive remarks about the stability of the quark matter. Above (below) the black dashed line, the 2+1-flavor (2-flavor) quark matter is more energetically stable, as indicated by the blue (orange) arrows. The two asterisks correspond to the two asterisks marked in the left panel.
@@
For a quark matter system to be stable, the matter pressure (comprising both the pQCD and CS contributions) must balance the bag pressure $-B$ at a baryon chemical potential $\mu_{\text{B}}\leq 930$ MeV  [^2] One could relax this condition by using the free nucleon mass $m_{p}\approx 938$ MeV instead, assuming a very long tunneling time to form a nucleus. This brings a negligible change to the range of the relevant parameter space., the energy per baryon of the iron element. We illustrate this with two benchmarks of $X=1.6$ and $2.2$ for both the 2+1- and 2-flavor systems by plotting their $p(\mu_{\text{B}},X)=p_{\rm pQCD}+p_{\rm CS}$ distributions in the left panel of Figure 2. The black dashed vertical line in the plot stands for $\mu_{\text{B}}=930$ MeV, and its intersections with the $p$ distributions imply the upper bounds on the bag parameter $B$ that allow stable quark matter for the different systems. As discussed in Ref. [10], the region with $p\simeq 0$ is also where $\alpha_{s}$ becomes large such that the pQCD results become unreliable. Depending on the choice of the $X$ parameter, either the 2+1- or 2-flavor quark matter could be more stable than the other one. To guide our discussion, we use the asterisks to mark the locations of $p^{1/4}$ at $\mu_{\text{B}}=930$ MeV for the $X=1.6$ benchmark of the 2+1-flavor scenario and the $X=2.2$ benchmark of the 2-flavor scenario.
@@
In the right panel of Figure 2, we show the consequent constraint plot on the $X$-$B^{1/4}$ plane that implies stable quark matter. The gray shaded region with $X\leq 1.42$ is the region that the pQCD calculations break down for $\mu_{\text{B}}\leq 930$ MeV (see Appendix A for details). In this region, no conclusive remarks based on pQCD can be made about the stability of the quark matter. For the region with $X>1.42$, the blue and orange lines denote the boundary of stable 2+1- and 2-flavor quark matter, respectively. For the regions between the blue solid and orange dashed lines, only the 2+1-flavor quark matter can be stable. The black dashed line shows the boundary of the relative stability between the 2+1-flavor and 2-flavor quark matter: the region above (below) the black dashed line indicates that the 2+1-flavor (2-flavor) quark matter is more energetically stable. Also note that there is some mild numerical uncertainty in the region around $X=1.42$, which does not change the general boundary feature. From this plot, one can see that for most $X$ values (except the uncertain region around $X=1.42$), the 2-flavor scenario is more stable when $B^{1/4}$ is small (below the black dashed line) and the other way around when $B^{1/4}$ is large (above the black dashed line). This is due to the fact that the strange flavor only becomes more populated when $\mu_{\text{B}}$ is high, thus contributing more to the pressure of the 2+1-flavor system, but will be suppressed in the low-$\mu_{\text{B}}$ regime where the strange quark mass matters more.

### v3

@@ Abstract
The determination of whether the ground state of baryon matter in Quantum Chromodynamics (QCD) is the ordinary nucleus or a quark matter state remains a long-standing question in physics. A critical parameter in this investigation is the bag parameter $B$, which quantifies the QCD vacuum energy and can be computed using nonperturbative methods such as Lattice QCD (LQCD). By combining the equation of state derived from perturbative QCD (pQCD) with the bag parameter to fit the LQCD-simulated data for isospin-dense matter, we address the stability of quark matter within the LQCD+pQCD framework. Our findings suggest that the current data imposes an upper bound on $B^{1/4}\lesssim 160$ MeV, approaching a conclusive statement on quark matter stability. Given the lower bound on $B$ from the quark condensate contribution to the vacuum energy, the stable 2-flavor quark matter remains possible, whereas the stable 2+1-flavor quark matter is excluded, assuming complete deconfinement and chiral-symmetry restoration and the reliability of pQCD at baryon chemical potentials around the proton mass. Additionally, we derive more general thermodynamic bounds on the quark matter energy-per-baryon and $B$, which, while weaker, provide complementary insights.
@@ 1 Introduction
It has long been thought that the global ground state of the baryon matter in quantum chromodynamics (QCD) should be the hadron state, while the quark matter state could only be accessible at environments of high temperature (such as the early universe or ion colliders) or high density (such as the core of a massive neutron star). However, the possible form of quark matter existence as a solitonic “bag” in the zero-pressure vacuum, also known as the “quark nugget”, has long ago been hypothesized and studied extensively by physicists such as Bodmer [1], Lee and collaborators [2, 3, 4, 5], and Witten [6], and still receives wide attention today. In contrast to the possible quark matter at the core of a neutron star, whose stability is supported by the external pressure, the stability of a quark nugget relies entirely on the energy configuration of the matter itself. Usually, a quark matter state is considered stable if its energy per baryon $\epsilon/n_{\text{B}}$ is less than around $930$ MeV, the averaged mass per nucleon of the iron element. So far whether this is true still remains an open question. The question of quark matter stability is of major importance since the implied stable quark nuggets could potentially extend the periodic table by introducing “exotic nuclei” with very large atomic and atomic mass numbers. Furthermore, stable quark nuggets, if produced in the early universe, could serve as a compelling dark matter candidate [6]. One such study of quark matter stability based on phenomenological models can be found in Ref. [7]. On the other hand, we aim to address this question using the fundamental QCD, which should be more robust than the known phenomenological models.
To investigate this question, one can start from the high-density limit, within which the thermodynamic equation of state (EOS) of the quark matter phase can be calculated based on perturbative QCD (pQCD) [8, 9, 10]. For instance, Ref. [11] has performed such calculations up to $\mathcal{O}(\alpha_{s}^{2})$ including the strange quark mass, which we mainly follow in this work. We note that partial results of the $\mathcal{O}(\alpha_{s}^{3})$ calculations are available in Refs. [12, 13], which we will discuss later in this study. The pressure from pQCD calculations is a positive function of the baryon chemical potential $\mu_{\text{B}}$. The additional formation of color superconducting (CS) phase from Bardeen-Cooper-Schrieffer condensates provides also positive but negligible pressure to the quark matter system in the perturbative regime. To achieve a quark matter state with zero external pressure, an additional negative contribution to the total pressure is required. This contribution could arise from the vacuum energy difference between the confined (chiral-symmetry breaking) and deconfined (chiral-symmetry restored) QCD phases, which is nonperturbative in nature.
@@
$\hskip 85.35826ptB\equiv-\frac{1}{4}\,\langle\Theta^{\mu}_{\;\mu}\rangle\penalty 10000\ ,\qquad\mbox{with}\quad\Theta^{\mu}_{\;\mu}=\frac{\beta(g)}{2g}G_{a}^{\mu\nu}G^{a}_{\mu\nu}+\sum_{q}\,m_{q}\,[1+\gamma_{m}(g)]\,\overline{q}q\penalty 10000\ .$ (1.1)
Here, $\Theta^{\mu}_{\;\mu}$ is the trace anomaly that depends on the renormalized fields and the QCD gauge coupling $g$ [16, 17]; $\beta(g)$ is the beta-function of the gauge coupling and $\gamma_{m}(g)$ is the quark mass anomalous dimension. The vacuum expectation value of the trace anomaly operator is defined at zero temperature and density in the confined vacuum. One way to obtain $B$ is to infer the values of the condensates from other observables (see for example Refs. [18, 19, 20, 21, 22, 23, 24]); however, this approach suffers from large systematic uncertainties, which we will discuss in more detail in Section 5. An alternative approach is to directly perform nonperturbative calculations of the thermodynamic properties of the baryon matter system, and lattice QCD (LQCD) is by far the most unique tool for this task.
In order to explore the transition from the hadron phase to the quark matter phase, one can either probe along the temperature direction or along the chemical potential direction (see for example Ref. [25] for a review on the QCD phase diagram). In this work, we focus on the latter and will comment on the former in Section 7. It is well known that simulating a system at finite $\mu_{\text{B}}$ (LQCDB) suffers from the fermion sign problem [26], making it challenging to extend the study to a high $\mu_{\text{B}}$. On the other hand, the isospin-dense matter system (LQCDI) is free of this problem (see Section 4), and thus can be studied up to a high isospin chemical potential $\mu_{\text{I}}$ to explore the deconfined quark matter phase. Recently, Refs. [27, 28] have performed LQCD simulations on a system of 6144 pions up to $\mu_{\text{I}}\sim 3.2$ GeV at zero temperature, in which they have identified the first-order phase transition to a Bose-Einstein condensate phase and a crossover to the CS phase [29, 30] through measuring the pressure $p$, energy density $\epsilon$, and speed of sound $c_{s}$ of the system. Based on the next-to-leading-order (NLO) calculations of the CS contribution to the EOS derived in Ref. [31], Ref. [32] has performed a matching between the LQCDI data and the pQCD+CS calculations to estimate the CS gap parameter. In this work, we take into account the $B$ parameter together with pQCD+CS to analyze the LQCDI data. This will not only provide a more complete description of the system, but also potentially examine the stability of quark matter.
@@ 2 Phases of baryon matter
$-\frac{\Omega}{V}=p=\mu_{\text{B}}\,n_{\text{B}}-\epsilon\penalty 10000\ .$ (2.1)
@@
Figure 1: Left panel: A schematic plot of the $p(\mu_{\text{B}})$ curves of the hadron matter phase and the quark matter phase at zero temperature, when there is only one intersection at $\mu_{\text{B}}=\mu_{\rm t}$ (green square point) between the two phases. The blue circle point at $\mu_{\text{B}}=\mu_{\rm h}$ with $p=0$ gives the energy per baryon of the (free) hadron matter as $\mu_{\rm h}$, while the red circle point at $\mu_{\text{B}}=\mu_{\rm q}$ gives that of the quark matter. The lower endpoint of the quark matter phase at $\mu_{\text{B}}=\mu_{0}$ and p=Bp=-B has $n_{\text{B}}=0$. The globally stable baryon matter phase (green solid line) follows the trajectory of the maximum $p$ at any given $\mu_{\text{B}}$. Right panel: The same as the left panel but with two intersections at $\mu_{\text{B}}=\mu_{{\rm t}_{1}}$ and $\mu_{{\rm t}_{2}}$ between the two phases.
@@
The ordinary hadron matter phase (blue dot-dashed line) starts from a nonzero $\mu_{\text{B}}=\mu_{\rm h}$ (bounded from below by the baryon mass), where both the pressure and the number density are zero. As for the quark matter phase (red dashed line), the starting chemical potential $\mu_{0}$ [^1] As we will discuss more in Section 3 and Appendix A, $\mu_{0}$ results mathematically from the breakdown of pQCD calculations when $\alpha_{s}$ becomes too large. While there is no rigorous proof, $\mu_{0}$ could also play a similar physical role for the quark matter as $\mu_{\rm h}$ for the hadron matter, the latter of which determines the minimum energy per baryon needed to produce a hadron state. could lie at a lower value, where $n_{\text{B}}$ is zero but the pressure is non-positive, p=Bp=-B with $B\geq 0$. As $\mu_{\text{B}}$ increases, there will be a zero-pressure point at $\mu_{\text{B}}=\mu_{\rm q}$ where the free quark matter state can exist with its energy per baryon given by $\epsilon/n_{\text{B}}=\mu_{\rm q}$ as derived from Eq. (2.1).
@@
The $p(\mu_{\text{B}})$ curves of the two phases are expected to have at least one intersection. In the high $\mu_{\text{B}}$ limit, one anticipates the quark matter phase (including the CS phase) to be the ultimate phase, and the hadron matter phase starting from $\mu_{\text{B}}=\mu_{\rm h}$ will eventually flow to the quark matter phase. In general, the $p(\mu_{\text{B}})$ curves of the two phases can intersect multiple times. For simplicity, we will consider the situation with only one intersection at $\mu_{\rm t}$ (shown in the left panel of Figure 1) and that with two at $\mu_{{\rm t}_{1,2}}$ (shown in the right panel of Figure 1). Using the Maxwell construction to patch the curves of the two different phase, [^2] Alternatively, the Gibbs construction further considers the electric charge chemical potential of the thermodynamic system. both the chemical potentials and pressures of the two phases should be continuous at the transition boundaries, i.e., $\mu_{1}=\mu_{2}=\mu_{\rm t}$ and $p_{1}(\mu_{\rm t})=p_{2}(\mu_{\rm t})$. For the case with one intersection in the left panel, one inevitably has $\mu_{\rm q}>\mu_{\rm h}$ or $\epsilon/n_{\text{B}}|_{\rm q}>\epsilon/n_{\text{B}}|_{\rm h}$, indicating an unstable quark matter state. For a fixed value of $\mu_{\rm t}$, $\mu_{\rm q}$ will increase along with increasing $B$, making the quark matter state more unstable. Conversely, $\mu_{\rm q}<\mu_{\rm h}$ for the case of two intersections in the right panel, resulting in a more stable quark matter state compared to the ordinary hadron matter state.
@@
Before concluding this section, we comment on the possible scenario that the hadron and quark matter phases are connected through a smooth crossover. In this case, one should expect $p(\mu_{\text{B}})=S(\mu_{\text{B}})p_{1}(\mu_{\text{B}})+[1-S(\mu_{\text{B}})]p_{2}(\mu_{\text{B}})$ with $S(\mu_{\text{B}})\in[0,1]$ as an infinitely differentiable switching function between the two phases [35, 36]. Since there is no clear boundary between the two phases, the simple picture of the quark matter pressure balancing the vacuum pressure only applies at asymptotically high $\mu_{\text{B}}$. In this study, we will assume that this picture is valid either due to the first-order nature of the phase transition or that the $\mu_{\text{B}}$ range of our interest is high enough to justify this assumption for a crossover phase transition.
@@ 3 Quark matter stability from pQCD calculations
At zero temperature and a high enough chemical potential, the grand potential or the pressure of the quark matter can be calculated perturbatively by treating the QCD coupling $\alpha_{s}$ as an expansion parameter [8, 9, 10, 11] and including the single quark loops, two-gluon irreducible vacuum diagrams, and the plasmon ring sum (see Ref. [11] for details). In this section, we follow Ref. [11] to calculate the thermodynamic properties of the quark matter system up to $\mathcal{O}(\alpha_{s}^{2})$. [^3] The plasmon ring sum contains terms of $\mathcal{O}(\alpha_{s}^{2}\ln\alpha_{s})$ resulting from the resummations, though for simplicity we still refer to the nominal order as $\mathcal{O}(\alpha_{s}^{2})$. Together with the unknown bag parameter $B$, which can only be calculated nonperturbatively, we derive the stability conditions for the quark matter.
@@
$X=\begin{cases}\dfrac{\mu_{\text{R}}}{\frac{2}{N_{c}}\left(\mu_{u}+\mu_{d}+\mu_{s}\right)}&\penalty 10000\ \text{[2+1-flavor]}\\[12.91663pt] \dfrac{\mu_{\text{R}}}{\frac{2}{N_{c}}\left(\mu_{u}+2\mu_{d}\right)}&\penalty 10000\ \text{[2-flavor]}\\[12.91663pt] \dfrac{\mu_{\text{R}}}{\mu_{u}-\mu_{d}}&\penalty 10000\ \text{[isospin-dense]}\end{cases}\penalty 10000\ ,$ (3.1)
@@
Figure 2: Left panel: The $p(\mu_{\text{B}},X)=p_{\rm pQCD}+p_{\rm CS}$ distributions of the benchmarks with $X=1.6$ and $2.2$ for the 2+1- and 2-flavor systems. The black dotted vertical line stands for $\mu_{\text{B}}=930$ MeV, whose intersections with the $p$ distributions imply the upper bounds on the bag parameter $B$ that allow stable quark matter. Right panel: The parameter region (white) of stable quark matter with trustable pQCD calculations. The regions above the blue solid and orange dashed lines exclude the existence of stable 2+1- and 2-flavor quark matter, respectively. The pQCD calculation breaks down for $\mu_{\text{B}}\leq 930$ MeV in the shaded gray region of $X\leq 1.42$ (see Appendix A for details) and thus cannot give conclusive remarks about the stability of the quark matter. Above (below) the black dashed line, the 2+1-flavor (2-flavor) quark matter is more energetically stable, as indicated by the blue (orange) arrows. The two asterisks correspond to the two asterisks marked in the left panel.
@@
For a quark matter system to be stable, the matter pressure (comprising both the pQCD and CS contributions) must balance the bag pressure B-B at a baryon chemical potential $\mu_{\text{B}}\leq 930$ MeV, [^4] One could relax this condition by using the free nucleon mass $m_{p}\approx 938$ MeV instead, assuming a very long tunneling time to form a nucleus. This brings a negligible change to the range of the relevant parameter space. the energy per baryon of the iron element. We illustrate this with two benchmarks of $X=1.6$ and $2.2$ for both the 2+1- and 2-flavor systems by plotting their $p(\mu_{\text{B}},X)=p_{\rm pQCD}+p_{\rm CS}$ distributions in the left panel of Figure 2. The black dotted vertical line in the plot stands for $\mu_{\text{B}}=930$ MeV, and its intersections with the $p$ distributions imply the upper bounds on the bag parameter $B$ that allow stable quark matter for the different systems. As discussed in Ref. [11], the region with $p\simeq 0$ is also where $\alpha_{s}$ becomes large such that the pQCD results become unreliable, which is manifested in that $n_{\text{B}}(\mu_{\text{B}}<\mu_{0})<0$ and $p(\mu_{\text{B}}=\mu_{0})=0$ (see Appendix A for details). Depending on the choice of the $X$ parameter, either the 2+1- or 2-flavor quark matter could be more stable than the other one. To guide our discussion, we use the asterisks to mark the locations of $p^{1/4}$ at $\mu_{\text{B}}=930$ MeV for the $X=1.6$ benchmark of the 2+1-flavor scenario and the $X=2.2$ benchmark of the 2-flavor scenario.
@@
In the right panel of Figure 2, we show the consequent constraint plot on the $X$-$B^{1/4}$ plane that implies stable quark matter. The gray shaded region with $X\leq 1.42$ is the region that the pQCD calculations break down for $\mu_{\text{B}}\leq 930$ MeV since $\mu_{0}\leq 930$ MeV. In this region, no conclusive remarks based on pQCD can be made about the stability of the quark matter. For the region with $X>1.42$, the blue and orange lines denote the boundary of stable 2+1- and 2-flavor quark matter, respectively. For the regions between the blue solid and orange dashed lines, only the 2+1-flavor quark matter can be stable. The black dashed line shows the boundary of the relative stability between the 2+1-flavor and 2-flavor quark matter: the region above (below) the black dashed line indicates that the 2+1-flavor (2-flavor) quark matter is more energetically stable. Also note that there is some mild numerical uncertainty in the region around $X=1.42$, which does not change the general boundary feature. From this plot, one can see that for most $X$ values (except the uncertain region around $X=1.42$), the 2-flavor scenario is more stable when $B^{1/4}$ is small (below the black dashed line) and the other way around when $B^{1/4}$ is large (above the black dashed line). This is due to the fact that the strange flavor only becomes more populated when $\mu_{\text{B}}$ is high, thus contributing more to the pressure of the 2+1-flavor system, but will be suppressed in the low-$\mu_{\text{B}}$ regime where the strange quark mass matters more.
Finally, we make a comment on the perturbative convergence of the calculations and point out a possible caveat of the stability analysis presented in this section, with the details given in Appendix A. In the $\mu_{\text{B}}$ range of our interest, we find that $\alpha_{s}(\mu_{\text{B}})\lesssim 0.6$ for $X\geq 1.47$ for both the 2- and 2+1-flavor scenarios. The fact that $(\frac{\alpha}{\pi})<1$ should indicate the applicability of pQCD subject to a complication of different $SU(3)_{C}$ group factors showing up at different loop orders. As we show in the Appendix using a simple massless three-flavor scenario with $X=2$, the $\mathcal{O}[(\frac{\alpha}{\pi})^{1}]$ and $\mathcal{O}[(\frac{\alpha}{\pi})^{2}]$ corrections, though being comparable in magnitude, are indeed subdominant compared to the $\mathcal{O}[(\frac{\alpha}{\pi})^{0}]$ contribution and of the same negative sign. Specifically, $\alpha_{s}(\mu_{\text{B}}=930\penalty 10000\ {\rm MeV})=0.42$, clearly satisfying the condition $(\frac{\alpha}{\pi})<1$. We further show in the Appendix that if the $\mathcal{O}[(\frac{\alpha}{\pi})^{3}]$ correction turns out to completely cancel the $\mathcal{O}[(\frac{\alpha}{\pi})^{2}]$ correction, i.e., leaving only the $\mathcal{O}[(\frac{\alpha}{\pi})^{1}]$ correction intact, our conclusion on quark matter stability still remains robust (see Section 5 for the conclusion based on the $(\frac{\alpha}{\pi})^{2}]$ calculations). Based on the partial results presented in Ref. [12], such a dramatic cancellation seems very unlikely to happen; moreover, the increasing trend of $\mu_{0}$ for high $X$ values suggests that higher-order corrections will further solidify our conclusion in this paper. However, until the complete calculations are done [13], there is no guarantee that the $\mathcal{O}[(\frac{\alpha}{\pi})^{3}]$ correction will not be larger than the sum of the $\mathcal{O}[(\frac{\alpha}{\pi})^{1}]$ and $\mathcal{O}[(\frac{\alpha}{\pi})^{2}]$ corrections with an opposite sign and invalidate the predictions made in this study.